A man jumps off a diving board into a pool. The equation h(t)=16T^2 + 40 gives the height,in feet, the man is above the pool t seconds after he jumps.
a) how high above the pool is the diver after the first second?

b) how high is the diving board?

Answer by Alan3354(69443)

A man jumps off a diving board into a pool. The equation h(t)=16T^2 + 40 gives the height,in feet, the man is above the pool t seconds after he jumps.
a) how high above the pool is the diver after the first second?
------------
h(t) = 16t^2 + 40
h(1) = 16 + 40 = 56 ft at 1 second
h(5) = 16*25 + 40 = 440 ft at 5 seconds
Do you see a problem with that?
------------
Use
h(t) = -16t^2 + 40
Find h(1)
====================
b) 40 feet